### v2

@@ 5 Experiments
In this section, we first introduce four video editing tasks that we focus on evaluating our AnyV2V framework: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation. We then assess the effectiveness of AnyV2V both quantitatively and qualitatively on these four tasks. For (1) prompt-based editing, we compare AnyV2V with SoTA text-guided video editing baselines. For the three novel tasks (2)(3)(4), as these tasks require reference images to guide video editing, current text-based video editing methods are incapable of performing such edits. We thus compare the performance of AnyV2V using different I2V generation backbones.
5.1 Tasks Definition
@@
5.2 Implementation Details
We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl. [62], ConsistI2V [44] and SEINE [12]. For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps. We use the DDIM [47] sampler and set $T$ to the default values of the selected I2V models. Following PnP [50], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections. During sampling, we apply text classifier-free guidance (CFG) [28] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits. We refer readers to the supplementary materials for more discussions on our implementation details and hyperparameter settings.
To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [6], style transfer model Neural Style Transfer (NST) [18], subject-driven image editing model AnyDoor [11], and identity-driven image editing model InstantID [51]. We experiment with only the successfully edited frames, which is crucial for our method.
5.3 Quantitative Evaluations
5.3.1 Prompt-based Editing.
@@ 5.3 Quantitative Evaluations
5.3.2 Reference-based Style Transfer; Identity Manipulation and Subject-driven Editing
@@
[…]
Task Prompt-based Editing
Method Human Evaluation $\uparrow$ CLIP Scores $\uparrow$
Alignment Overall CLIP-Text CLIP-Image
Tune-A-Video [55] 15.2% 2.1% 0.2902 0.9704
[…]
AnyV2V (I2VGen-XL) 69.7% 46.2% 0.2932 0.9652
@@
5.4 Qualitative Results
5.4.1 Prompt-based Editing.
By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing. In our experiment, we used InstructPix2Pix [6] for the first frame edit. Shown in Figure 3, our method correctly places a party hat on an old man’s head and successfully turns the color of an airplane to blue, while preserving the background and keeping the fidelity to the source video. Comparing our work with the three baseline models TokenFlow [20], FLATTEN [15], and Tune-A-Video [55], the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt. The color tone and object shapes are also tilted. It is also worth mentioning that our approach is far more consistent on some motion tasks such as adding snowing weather, due to the I2V model’s inherent support for animating still scenes. The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modelling.
Figure 3: AnyV2V is robust in a wide range of prompt-based editing tasks while maintaining the background. The generated results align the most with the text prompt and also maintain high motion consistency.
5.4.2 Reference-based Style Transfer.
5.4.3 Subject-driven Editing.
5.4.4 Identity Manipulation.
Model CLIP-Image $\uparrow$
AnyV2V (I2VGen-XL) 0.9648
AnyV2V (I2VGen-XL) w/o T. Injection 0.9652
AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection 0.9637
5.5 Ablation Study
To verify the effectiveness of our design choices, we conduct an ablation study by iteratively disabling the three core components in our model: temporal feature injection, spatial feature injection, and DDIM inverted latent as initial noise. We use AnyV2V (I2VGen-XL) and a subset of 20 samples in this ablation study and report both the frame-wise consistency results using CLIP-Image score in Table 3 and qualitative comparisons in Figure 5. We provide more ablation analysis of other design considerations of our model in the supplementary materials.
5.5.1 Effectiveness of Temporal Feature Injection.
[…]
5.5.2 Effectiveness of Spatial Feature Injection.
As shown in Table 3, we observe a drop in the CLIP-Image score after removing the spatial feature injection mechanisms from our model, indicating that the edited videos are not smoothly progressed across consecutive frames and contain more appearance and motion inconsistencies. Further illustrated in the third row of Figure 5, removing spatial feature injection will often result in incorrect subject appearance and pose (as shown in the “ballet dancing” case) and degenerated background appearance (evident in the “couple sitting” case). These observations demonstrate that directly generating edited videos from the DDIM inverted noise is often not enough to fully preserve the source video structures, and the spatial feature injection mechanisms are crucial for achieving better editing results.
5.5.3 DDIM Inverted Noise as Structural Guidance.
Finally, we observe a further decrease in CLIP-Image scores and a significantly degraded visual appearance in both examples in Figure 5 after replacing the initial DDIM inverted noise with random noise during sampling. This indicates that the I2V generation models become less capable of animating the input image when the editing prompt is completely out-of-domain and highlights the importance of the DDIM inverted noise as the structural guidance of the edited videos.
6 Conclusion
In this paper, we present AnyV2V, a new unified framework for video editing. Our framework is training-free, highly cost-effective, and can be applied to any I2V model. To perform video editing with high precision, we propose a two-stage approach to first edit the initial frame of the source video and then condition an image-to-video (I2V) model with the initial frame. Comprehensive experiments have shown that our method achieves outstanding outcomes across a broad spectrum of applications and shows a high level of controllability.
7 Limitations
7.0.1 Inaccurate Edit from Image Editing Models.
@@ 7 Limitations
As our method relies on an initial frame edit, the image editing models are used. However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [32]. For example, in the subject-driven video editing task, we found that AnyDoor [11] requires several tries to get a good editing result. Efforts are required in manually picking a good edited frame. We expect that in the future better image editing models will minimize such effort.
7.0.2 Limited ability of I2V models.
@@
We found that the results from our method cannot follow the source video motion if the motion is fast (e.g. billiard balls hitting each other at full speed) or complex (e.g. a person clipping her hair). One possible reason is that the current popular I2V models are generally trained on slow-motion videos, such that lacking the ability to regenerate fast or complex motion even with motion guidance. We anticipate that the presence of a robust I2V model can address this issue.
8 Potential Negative Impacts
@@
Supplementary Material
Appendix 0.A Discussion on Model Implementation Details
0.A.1 U-Net Layers for Feature Injection
0.A.2 Ablation Analysis on Feature Injection Thresholds
0.A.2.1 Effect of Spatial Injection Thresholds $\tau_{conv},\tau_{sa}$
Source
$\tau_{c,s}=0$
$\tau_{c,s}=0.2T$
$\tau_{c,s}=0.5T$
$\tau_{c,s}=0.7T$
$\tau_{c,s}=T$
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/source/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/0/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/2/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/5/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/7/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/10/015
$\tau_{c,s}=0$
$\tau_{c,s}=0.2T$
$\tau_{c,s}=0.5T$
$\tau_{c,s}=0.7T$
$\tau_{c,s}=T$
Figure 7: Hyperparameter study on spatial feature injection. We find that $\tau_{sa}=0.2T$ is the best setting for maintaining the layout and structure in the edited video while not introducing unnecessary visual details from the source video. $\tau_{c,s}$ represents $\tau_{conv}$ and $\tau_{sa}$. (Editing prompt: teddy bear running. Videos clickable in Adobe Acrobat.)
0.A.2.2 Effect of Temporal Injection Threshold $\tau_{ta}$
Source
$\tau_{ta}=0$
$\tau_{ta}=0.2T$
$\tau_{ta}=0.5T$
$\tau_{ta}=0.7T$
$\tau_{ta}=T$
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/src/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.2_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.5_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.7_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp1.0_stp50/frames/frame_015
$\tau_{ta}=0$
$\tau_{ta}=0.2T$
$\tau_{ta}=0.5T$
$\tau_{ta}=0.7T$
$\tau_{ta}=T$
Figure 8: Hyperparameter study on temporal feature injection. We find that $\tau_{ta}=0.5T$ to be the optimal setting as it balances motion alignment, motion consistency, and video fidelity. (Editing prompt: darth vader walking. Videos clickable in Adobe Acrobat.)
0.A.3 Backbones Performance
Among the three backbones of our framework, We found that AnyV2V (I2VGen-XL) tends to be the most robust model both qualitatively and quantitatively. It has a good generalization ability to produce consistent motions in the video with high visual quality. AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its data used during the training stage, thus harming the visual quality. AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking. One interesting property of our framework is that although in some cases the motion does not fully align with the source video, it produces video with high motion consistency and reasonable motion that is highly similar to the source video, indicating the flexibility of how the video can be edited while maintaining the video fidelity.
Appendix 0.B Human Evaluation Dataset
Our human evaluation dataset contains in total of 89 samples. For prompt-based editing, we employed InstructPix2Pix [6] to compose the examples. Topics include swapping objects, adding objects, and removing objects. For subject-driven editing, we employed AnyDoor [11] to replace objects with reference subjects. For Neural Style Transfer, we employed NST [18] to compose the examples. For identity manipulation, we employed InstantID [51] to compose the examples. See Table 4 for the statistic.
Table 4: Number of entries for Video Editing Evaluation Dataset
Category Number of Entries Image Editing Models Used
Prompt-based Editing 45 InstructPix2Pix [6]
Reference-based Style Transfer 20 NST [18]
Identity Manipulation 13 InstantID [51]
Subject-driven Editing 11 AnyDoor [11]
Total 89

### v3

@@ 4 Experiments
Figure 3: AnyV2V is robust in a wide range of prompt-based editing tasks while preserving the background. The results align the most with the text prompt and maintain high motion consistency.
4.1 Implementation Details
We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl. [59], ConsistI2V [42] and SEINE [10]. For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps. We use the DDIM [44] sampler and set $T$ to the default values of the selected I2V models. Following PnP [47], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections. During sampling, we apply text classifier-free guidance (CFG) [22] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits. To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [4], style transfer model Neural Style Transfer (NST) [15], subject-driven image editing model AnyDoor [9], and identity-driven image editing model InstantID [49]. We experiment with only the successfully edited frames, which is crucial for our method. We conducted all the experiments on a single Nvidia A6000 GPU. To edit a 16-frame video, it requires around 15G GPU memory and around 100 seconds for the whole inference process. We refer readers to Appendix A for more discussions on our implementation details and hyperparameter settings.
[…]
Method Human Evaluation $\uparrow$ CLIP Scores $\uparrow$
Alignment Overall CLIP-Text CLIP-Image
Tune-A-Video [53] 15.2% 2.1% 0.2902 0.9704
[…]
AnyV2V (I2VGen-XL) 69.7% 46.2% 0.2932 0.9652
4.2 Evaluation Results
As shown in Figure 3 and 4, AnyV2V can perform the following video editing tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation. we compare AnyV2V against three baseline models Tune-A-Video [53], TokenFlow [16] and FLATTEN [12] for (1) prompt-based editing. Since there exists no publicly available baseline method for task (2) (3) (4), we evaluate the performance of three I2V generation models under AnyV2V. We included a more comprehensive evaluation in Appendix B.
Prompt-based Editing Unlike the baseline methods [53, 16, 12] that often introduce unwarranted changes not specified in the text commands, AnyV2V utilizes the precision of image editing models to ensure only the targeted areas of the scene are altered, leaving the rest unchanged. Combining AnyV2V with the instruction-guided image editing model InstructPix2Pix [4], AnyV2V accurately places a party hat on an elderly man’s head and correctly paints the plane in blue. Additionally, it maintains the original video’s background and fidelity, whereas, in comparison, baseline methods often alter the color tone and shape of objects, as illustrated in Figure 3. Also, for motion tasks such as adding snowing weather, I2V models from AnyV2V provide inherent support for animating the snowing while the baseline methods would result in flickering. For quantitative evaluations, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting, and also compute the metrics CLIP-Text for text alignment and CLIP-Image for temporal consistency. In detail, CLIP-Text is computed by the average cosine similarity between text embeddings from CLIP model [41] and CLIP-Image is computed in the same way but between image embeddings for every pair of consecutive frames. Table 2 shows that AnyV2V generally achieves high text-alignment and temporal consistency, while AnyV2V with I2VGen-XL backbone is the most preferred method because it does not edit the video precisely.
Style Transfer, Subject-Driven Editing and Identity Manipulation For these novel tasks, we stress the alignment with the reference image instead of the text prompt. As shown in Figure 4, we can observe that for task (2), AnyV2V can capture one particular style that is tailor-made, and even if the style is not learned by the text encoder. In the examples, AnyV2V captures the style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately. For task (3), AnyV2V can replace the subject in the video with other subjects even if the new subject differs from the original subject. In the examples, a cat is replaced with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video. A car is replaced by our desired car while the wheel is still spinning in the edited video. For task (4), AnyV2V can swap a person’s identity to anyone. We report both human evaluation results and find that AnyV2V with I2VGen-XL backbone is the most preferred method in terms of reference alignment and overall performance, plotted in Table 5, which is in Appendix B.
I2V Backbones We find that AnyV2V (I2VGen-XL) tends to be the most robust both qualitatively and quantitatively. It has a good generalization ability to produce consistent motions in the video with high visual quality. AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its training data, thus harming the visual quality. AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking.
Editing Video beyond Training Frames of I2V model Current state-of-the-art I2V models [10, 42, 59] are mostly trained on video data that contains only 16 frames. To edit videos that have length beyond the training frames of the I2V model, an intuitive approach would be generating videos in an auto-regressive manner as used in ConsistI2V [42] and SEINE [10]. However, we find that such an experiment setup cannot maintain semantic consistency in our case. As many works in extending video generation exploit the initial latent to generate longer video [40, 54], we leverage the longer inverted latent as the initial latent and force an I2V model to generate longer frames of output. Our experiments found that the inverted latent contains enough temporal and semantic information to allow the generated video to maintain temporal and semantic consistency, as shown in Figure 5.
Figure 5: AnyV2V can edit video length beyond the training frame while maintaining motion consistency. The first row is the source video frames while the second rows are the edited. The editing prompt of the image was “turn woman into a robot” using image model InstructPix2Pix [4].
4.3 Ablation Study
To verify the effectiveness of our design choices, we conduct an ablation study by iteratively disabling the three core components in our model: temporal feature injection, spatial feature injection, and DDIM inverted latent as initial noise. We use AnyV2V (I2VGen-XL) and a subset of 20 samples in this ablation study and report both the frame-wise consistency results using CLIP-Image score in Table 3 and qualitative comparisons in Figure 6. We provide more ablation analysis of other design considerations of our model in the Appendix.
[…]
Effectiveness of Spatial Feature Injection As shown in Table 3, we observe a drop in the CLIP-Image score after removing the spatial feature injection mechanisms from our model, indicating that the edited videos are not smoothly progressed across consecutive frames and contain more appearance and motion inconsistencies. Further illustrated in the third row of Figure 6, removing spatial feature injection will often result in incorrect subject appearance and pose (as shown in the “ballet dancing” case) and degenerated background appearance (evident in the “couple sitting” case). These observations demonstrate that directly generating edited videos from the DDIM inverted noise is often not enough to fully preserve the source video structures, and the spatial feature injection mechanisms are crucial for achieving better editing results.
DDIM Inverted Noise as Structural Guidance Finally, we observe a further decrease in CLIP-Image scores and a significantly degraded visual appearance in both examples in Figure 6 after replacing the initial DDIM inverted noise with random noise during sampling. This indicates that the I2V generation models become less capable of animating the input image when the editing prompt is completely out-of-domain and highlights the importance of the DDIM inverted noise as the structural guidance of the edited videos.
Model CLIP-Image $\uparrow$
AnyV2V (I2VGen-XL) 0.9648
AnyV2V (I2VGen-XL) w/o T. Injection 0.9652
AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection 0.9637
5 Conclusion
In this paper, we presented AnyV2V, a novel unified framework for video editing. Our framework is training-free, highly cost-effective, and can be applied to any image editing model and I2V generation model. To perform video editing with high precision, we propose a two-stage approach to first edit the initial frame of the source video and then condition an I2V model with the edited first frame and the source video features and inverted latents to produce the edited video at any length. Comprehensive experiments have shown that our method achieves outstanding outcomes across a broad spectrum of applications that are beyond the scope of existing SOTA methods while achieving superior results on both common video metrics and human evaluation. For future work, we aim to find a tuning-free method to bridge the I2V properties into T2V models so that we can leverage existing strong T2V models for video editing.
Appendix
Appendix A Discussion on Model Implementation Details
A.1 U-Net Layers for Feature Injection
A.2 Ablation Analysis on Feature Injection Thresholds
Effect of Spatial Injection Thresholds $\tau_{conv},\tau_{sa}$
Figure 8: Hyperparameter study on spatial feature injection. We find that $\tau_{sa}=0.2T$ is the best setting for maintaining the layout and structure in the edited video while not introducing unnecessary visual details from the source video. $\tau_{c,s}$ represents $\tau_{conv}$ and $\tau_{sa}$. (Editing prompt: teddy bear running. The experiment was conducted with the I2Vgen-xl backbone.
Effect of Temporal Injection Threshold $\tau_{ta}$
Figure 9: Hyperparameter study on temporal feature injection. We find that $\tau_{ta}=0.5T$ to be the optimal setting as it balances motion alignment, motion consistency, and video fidelity. (Editing prompt: darth vader walking. The experiment was conducted with the SEINE backbone.
Appendix B Evaluation
B.1 Tasks Definition
@@
B.2 Quantitative Evaluations
Prompt-based Editing
@@ B.2 Quantitative Evaluations
Reference-based Style Transfer; Identity Manipulation and Subject-driven Editing
@@
B.3 Qualitative Results
Prompt-based Editing
By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing. In our experiment, we used InstructPix2Pix [4] for the first frame edit. Shown in Figure 3, our method correctly places a party hat on an old man’s head and successfully turns the color of an airplane to blue, while preserving the background and keeping the fidelity to the source video. Comparing our work with the three baseline models TokenFlow [16], FLATTEN [12], and Tune-A-Video [53], the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt. The color tone and object shapes are also tilted. It is also worth mentioning that our approach is far more consistent on some motion tasks such as adding snowing weather, due to the I2V model’s inherent support for animating still scenes. The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modeling.
Reference-based Style Transfer
Subject-driven Editing
Identity Manipulation
B.4 Human Evaluation
B.4.1 Dataset
Our human evaluation dataset contains a total of 89 samples that have been collected from https://www.pexels.com. For prompt-based editing, we employed InstructPix2Pix [4] to compose the examples. Topics include swapping objects, adding objects, and removing objects. For subject-driven editing, we employed AnyDoor [9] to replace objects with reference subjects. For Neural Style Transfer, we employed NST [15] to compose the examples. For identity manipulation, we employed InstantID [49] to compose the examples. See Table 4 for the statistic.
Table 4: Number of entries for Video Editing Evaluation Dataset
Category Number of Entries Image Editing Models Used
Prompt-based Editing 45 InstructPix2Pix [4]
Reference-based Style Transfer 20 NST [15]
Identity Manipulation 13 InstantID [49]
Subject-driven Editing 11 AnyDoor [9]
Total 89
B.4.2 Interface
In the evaluation setup, we provide the evaluator with generated images from both the baseline models and the AnyV2V models. Evaluators are tasked with selecting videos that best align with the provided prompt or the reference image. Additionally, they are asked to express their overall preference for the edited videos. For a detailed view of the interface used in this process, please see Figure 10.
Figure 10: The interface of individual evaluation.
Appendix C Preliminary
C.1 Image-to-Video (I2V) Generation Models
$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (3)
C.2 DDIM Inversion
C.3 Plug-and-Play (PnP) Diffusion Features
@@
Appendix D Discussion
D.1 Limitations
@@
Inaccurate Edit from Image Editing Models. As our method relies on an initial frame edit, the image editing models are used. However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [28]. For example, in the subject-driven video editing task, we found that AnyDoor [9] requires several tries to get a good editing result. Efforts are required in manually picking a good edited frame. We expect that in the future better image editing models will minimize such effort.
@@
Limited ability of I2V models. We found that the results from our method cannot follow the source video motion if the motion is fast (e.g. billiard balls hitting each other at full speed) or complex (e.g. a person clipping her hair). One possible reason is that the current popular I2V models are generally trained on slow-motion videos, such that lacking the ability to regenerate fast or complex motion even with motion guidance. We anticipate that the presence of a robust I2V model can address this issue.
D.2 License of Assets
For image editing models, InstructPix2Pix [4] inherits Creative ML OpenRAIL-M License as it is built upon Stable Diffusion. Neural Style Transfer [15] is under Creative Commons Attribution 4.0 License, and code samples are licensed under the Apache 2.0 License. InstantID [49] is under Apache License 2.0. AnyDoor [9] is under the MIT License.
For baselines, Tune-A-Video [53] is under Apache License 2.0, TokenFlow [16] is under MIT License andFLATTEN [12] is under Apache License 2.0.
For the human evaluation dataset, the dataset has been collected from https://www.pexels.com, with all data governed by the terms outlined at https://www.pexels.com/license/.
We decide to release AnyV2V code under the Creative Commons Attribution 4.0 License for easy access in the research community.
D.3 Societal Impacts
Postive Social Impact. AnyV2V has the potential to significantly enhance the capabilities of video editing systems, making it easier for a wider range of users to manipulate images. This could have numerous positive social impacts, as users would be able to achieve their editing goals without needing professional editing knowledge, such as using Photoshop or painting.
@@ D.3 Societal Impacts
D.4 Safeguards
It is crucial to implement proper safeguards and responsible AI frameworks when developing user-friendly video editing systems. For the human evaluation dataset, we manually collect a diverse range of images to ensure a balanced representation of objects from various domains. We only collect images that are considered safe.
D.5 Ethical Concerns for Human Evaluation
We believe our proposed human evaluation does not incur ethical concerns due to the following reasons: (1) the study does not involve any form of intervention or interaction that could affect the participants’ well-being. (2) Rating video content does not involve any physical or psychological risk, nor does it expose participants to sensitive or distressing material. (3) The data collected from participants will be entirely anonymous and will not contain any identifiable private information. (4) Participation in the study is entirely voluntary, and participants can withdraw at any time without any consequences.
D.6 New Assets
Our paper introduces several new assets including a human evaluation dataset and demo videos generated by AnyV2V. Each asset is thoroughly documented, detailing its creation, usage, and any relevant methodologies. The human evaluation dataset documentation includes details on how participant consent was obtained. The demo videos are provided as an anonymized zip file to comply with submission guidelines, with detailed instructions for use. All assets are shared under an open license to facilitate reuse and further research.
Appendix E More AnyV2V Showcases
Figure 11: AnyV2V becomes an instruction-based video editing tool when plugged with instruction-guided image editing models like InstructPix2Pix [4]. Prompt used “Turn the couple into robots”, “Turn horse into zebra”, and “Turn the sand into snow”.
Figure 12: The more recent model InstantStyle [48] can seamlessly plug in AnyV2V to perform reference-based style transfer video editing. The reference art image in the bottom left corner is used to retrieve the first image edit.
Figure 13: AnyV2V can perform subject-driven video editing with a single image reference, using a subject-driven image editing model like AnyDoor [9]. The subject image in the bottom left corner is used to retrieve the first image edit.